OpenStudy (girlgoyleh.):

What is the value of the expression 3 with the exponent of -3

OpenStudy (anonymous):

So it's 3^-3?

OpenStudy (girlgoyleh.):

yes

OpenStudy (campbell_st):

the negative exponent indicates a fraction so its \[\frac{1}{3^3}\] you can finish it off by finding the value of 3^3

OpenStudy (anonymous):

That's like saying -3 x -3 x -3 -3 x -3 = 9 x -3 = -27 Do you understand?

OpenStudy (girlgoyleh.):

yes, thank you guys, i just wanted to make sure it was a fraction. does that mean that every question like that will be a fraction too?

OpenStudy (campbell_st):

well if its written in the same form it will be

OpenStudy (girlgoyleh.):

1/-27 isnt one of the answer choices

OpenStudy (anonymous):

The answer is -27. It's not a fraction.

OpenStudy (girlgoyleh.):

-27 isnt one of the answer choices

OpenStudy (anonymous):

Then we're not understand the problem correctly. Try writing it with the equation feature on openstudy.

OpenStudy (girlgoyleh.):

its -81 right?

OpenStudy (girlgoyleh.):

\[3^{-4}\]

OpenStudy (anonymous):

That's 81. Positive 81.

OpenStudy (anonymous):

-3 * -3 * - 3 * - 3 = 81

OpenStudy (girlgoyleh.):

thanks,